# Fast and Robust Flocking of Protesters on Street Networks

### v1

@@ Abstract
We propose a simple model of protesters scattered throughout a city who want to gather into large and mobile groups. This model relies on random walkers on a street network that follow tactics built from a set of basic rules. Our goal is to identify the most important rules for fast and robust flocking of walkers. We explore a wide set of tactics and show the central importance of a specific rule based on alignment. Other rules alone perform poorly, but our experiments show that combining alignment with them enhances flocking, and that obtained groups are then remarkably robust.
Keywords : street networks, gathering, flocking, tactic design, robustness, protests
@@ 1 Introduction
To do so, we model the city as a network of streets and intersections in Section 2. We then assume that protesters are biased random walkers on such a network. In Section 3, we present the set of basic rules that compose tactics walkers can use to move. We define, in Section 4, what flocking features groups must achieve and run extensive experiments to measure how effective each tactic is regarding those. We also explore the robustness of groups, measuring how they reform if adversary forces break them up while following an effective tactic. In Section 5, we compare our approach to related work on flocking and on formal analysis of protests. We emphasise the absence of a protocol that can solve our problem. Finally, we summarise our contributions in Section 6.
@@ 2.2 Walkers
Given a network $G=(V,E)$, we consider a set $W$ of walkers numbered from $1$ to $|W|$. We denote the location of walker $i$ at time $t$ by $x_{i}(t)=v$, with $v\in V$. At each time step $t$, walker $i$ moves to node $x_{i}(t+1)\in N(v)$. Rules we present in Section 3 determine the choice of $x_{i}(t+1)$.
For any link $(u,v)$ in $E$, we also define the flux of walkers from $u$ to $v$ as $J_{u\rightarrow v}(t)=|\{i,x_{i}(t-1)=u,x_{i}(t)=v\}|$.
We call group the set of walkers at a given node $v$ at a given time $t$: $g_{v}(t)=\{i,x_{i}(t)=v\}$. We denote by $n_{v}(t)=|g_{v}(t)|$ the number of walkers located at node $v$ at time $t$. We denote by $g(t)=|\{g_{v}(t),v\in V,g_{v}(t)\neq\emptyset\}|$ the number of non-empty groups at step $t$.
2.3 Discretization
The links of a street network generally represent street segments of very heterogeneous lengths [3]. Then, moves from a node to another one may have very different duration. In order to model this, we use a classical discretization procedure [4] that we describe below. It consists in splitting each link of the street network into pieces connected by evenly spaced nodes, see Figure 2.
More formally, given a link $(u,v)$ representing a piece of street of length $L$ in a street network and a discretization step $\delta$, we split each undirected link $(u,v)$ into $k$ new links, with $k=\left\lceil\frac{L}{\delta}\right\rceil$ as follows. Let us denote a set of new nodes $V(u,v)=\{v_{0},v_{1},...,v_{k}\}$ with the nodes from the original network being $v_{0}=u$ and $v_{k}=v$. We then define a set of new links $E(u,v)=\{(w_{i},w_{i+1}),i=0,1,...,k\}$. Then we define the set of all nodes $V=\bigcup_{u,v}V(u,v)$ and the set of links $E=\bigcup_{u,v}E(u,v)$. Finally, we model the street network by the graph $G=(V,E)$.
Figure 2: A piece of the discretized street network around Place de la Nation in Paris.
In the obtained graph, each link represents a street slice of length close to $\delta$. Then, the walkers defined in previous section consistently make a move of length approximately $\delta$ at each time step.
@@ 3.1 Available information
We assume that walkers have very limited capabilities: they have no access to the actual location of other walkers, even on neighbor nodes, and they have no long-term memory and no communication protocol. Instead, we assume they only have access to estimates of aggregated observables, such as the number of walkers on neighbor nodes and the flow intensity on surrounding links, and a short-term memory.
@@ 3.2 Criteria
  • Propulsion. The walker never goes back to its previous location $u$, and otherwise makes random moves: the criterion has value 00 for $u$ and value $1$ for other neighbors of $v$.
@@ 3.3 Walking rules
This logit equation is widely used in the literature because it has many convenient features [5]. In particular, it ensures $\omega$ is a positive and monotonic function of the criterion: the greater $C_{u,v,w}(t)$, the greater the chances to move to node $w$ at time $t$. It is also compatible with criteria having both positive and negative values, like the flux-based one below.
We conducted experiments with other functions with similar properties to equation 1. They displayed no significant differences.
@@ 3.4 Tactics
where $\alpha_{C}$ is the coefficient of criterion $C$, with $\sum_{C\in\mathcal{C}}\alpha_{C}=1$. Therefore, a tactic is defined by a set of criteria and their coefficients.
We call strict tactic one that has all its coefficients, except one, set to zero and then always follows the same criterion.
@@
Notice that we may see a tactic as an inhomogeneous Markov chain, where transition probabilities are time-dependent. Still, at any given time, each walker makes a transition from its current state $(u,v)$ giving its previous location $u$ and its current location $v$, to its next state $(v,w)$ giving its current location $v$ and its next location $w$.
@@ 3.5 Baseline
More formally, for each node $v$ at time $t$, a unique random neighbor $u$ of $v$ is chosen and all walkers $i$ such that $x_{i}(t)=v$ move to $u=x_{i}(t+1)$. This is equivalent to purely random walks of groups until they meet another group.
@@ 4.1 Flocking metrics
Definition 4.1 (cluster, gathering score $\rho(t)$).
@@
Definition 4.2 (mobility score $\mu(t)$).
The mobility $\mu_{i}(t)$ of walker $i$ is the number of distinct nodes a walker has already visited at time $t$: $\mu_{i}(t)=|\{v,\exists t^{\prime}\leq t,x_{i}(t^{\prime})=v\}|$. The mobility score $\mu(t)$ is the average walker mobility: $\mu(t)=\frac{1}{|W|}\sum_{i\in W}\mu_{i}(t)$.
@@
Definition 4.3 (sprawling score $\sigma(t)$).
@@
4.2 Behaviors on a single street
Our networks are essentially composed of streets, modeled as lines of nodes by the discretization presented in Section 2.3 (Figure 2). This plays an important role in walker dynamics, therefore we investigate this special case in this Section.
Figure 3 shows space-time diagrams of three different strict tactics, i.e. the ones using only one rule, for $100$ walkers on a line of $100$ nodes. These space-time diagrams display the evolution of walker locations over time. The horizontal axis represents the nodes, numbered from $1$ to $100$. The vertical axis represents time steps, numbered from $1$ to $200$. Then, the color at coordinates $(x,y)$ represents the number of walkers at node $x$ after $y$ time steps, namely $n_{x}(y)$. The color is white if this number is 00, and it is the darkest if this number is the maximal value for the considered run. Walkers are initially uniformly distributed on all nodes.
Figure 3: Space-time diagrams of three tactics on a line using only (from left to right) the Random, Attraction, and Alignment rule. The gathering score is expressed relatively to the maximum number of walkers in a given run.
With the Random and with the Attraction strict tactic (Figure 3, left and middle), the largest obtained groups during the run both always contain less than 10 walkers (respectively 6 and 8 on the Figure) which never makes up to 5% of all walkers. However, the two tactics have distinct behaviors: while the Random one leads to walker constantly moving between groups, the Attraction one leads to stable groups. These groups fail to move along the line, though.
With Alignment (Figure 3, right), a more interesting pattern emerges: all walkers gather into a unique cluster that achieves flocking since it moves all along the line.
In conclusion, the Alignment strict tactic is sufficient to obtain flocking on a street. However, the groups we obtain eventually reach the street extremities, which raise the question of their behavior at street intersections.
@@ 4.3 Extensive exploration of tactics
  • we consider $N$ walkers initially distributed uniformly at random in the network,
@@
  • we set $\beta$ large enough to ensure that each walking rule strictly follows its criterion,
@@
All strict tactics have very poor scores, except the Alignment one. This is a consequence from our previous Section results, where it was the only to generate large and mobile groups. Other strict tactics failed to achieve those patterns in a street, and therefore fail on an entire city.
Attraction, Follow and Random strict tactic have almost no dynamic as walkers do not explore much the network and do not form large groups. Propulsion however displays an excellent mobility score with an almost linear growth on Figure 5, as walkers never go back to their previous location.
Figure 5: Plots showing the evolution of gathering and mobility scores for the strict tactics and the best tactic. Gathering score is in log-log scale.
@@ 4.4 Best tactics
Finally, Figure 6 displays the sprawling score for all considered tactics. We observe that the best tactic produces a greater sprawling score than Alignment. The sprawling, for those two tactics, is due to groups following each others when they detect another group on a neighbor node, without necessarily merging with it.
@@
Figure 6: Plot showing the evolution of the sprawling scores for the strict tactics and the best tactic. Legend is the same as in Figure 5.
@@
However, with a larger sprawling, a cluster keeps most of its groups unchanged after an intersection. Indeed, the first of its groups to reach the intersection node might split into multiple groups, as described above. However, if using an Alignment-based tactic, next walkers will then all follow the biggest of the new groups that came from the split. This shows an effective tactic can benefit from a little bit of sprawling to maintain its clusters size.
@@ 4.5 Interpretation
However, this rule alone is not perfectly effective at avoiding splitting. The walkers in the first group in a cluster will not always all choose the same node at an intersection, as illustrated in the two bottom pictures of Figure 7, as the different possible nodes all have a flux equal to 00.
In this context, the Attraction rule allows walkers that split in the least chosen direction at an intersection to go backward and avoid loosing sight of the cluster.
@@
Figure 7 also shows why the baseline is not good enough: walkers do not move forward at every step. However we could obtain this with a collective propulsion (a collective decision with the Propulsion rule instead of the Random one), such tactic neither guaranty groups merge when they cross path. That is the strength of Alignment rule: walkers merge not only if they fall on the same cell, but also if they simply exchange their position.
4.6 Group robustness
It is crucial for protesters to form groups that resist adversary forces that may break them up. In order to explore this robustness, we model break ups as walkers suddenly following the Random strict tactic for one step. In this way, a group located at a given node splits into smaller groups that move to neighbor nodes, in a way similar to a group of protesters targeted by adversary forces.
More formally, we perform the following experiment: we run the Alignment strict tactic for $100$ steps, then we run the Random strict tactic for one step, and we finally run the Alignment strict tactic again. We perform the same experiment with the best tactic (Section 4.5) instead of the strict Alignment one.
Figure 8 displays the average observed scores over ten runs of each experiment, together with results of the same experiments without the Random strict tactic step. Since the mobility score is trivially increased by these experiments, which brings no useful insight, we do not display this metric.
Figure 8: Plots of robustness experiments, similar to Figure 5. We display the sprawling and gathering scores for the Alignment strict tactic and the best tactic, as well as for the experiments where all walkers perform a random move at time $100$.
The break up simulation clearly appears on both plots. The gathering score plots show that large groups reappear shortly after break ups. The two tactics do not have the same behaviour, though. The strict Alignment tactic needs some time to reach gathering scores close to its pre-break up ones (green curve in Figure 8). Instead, the best tactic is strikingly efficient in quickly recovering from break ups. Indeed, large groups do not only reappear shortly after the break up, they do so even faster than before. This brings its gathering score (black curve in Figure 8) to quickly catch up with the case with no break up (purple curve).
We also observe that, before going back to normal, the sprawling score experiences a sharp increase immediately followed by a sharp decrease. The increase is due to the fact that groups in each cluster first split into several neighbor groups, thus forming a larger cluster. Then, the new groups tend to go forward, due to the Alignment rule, and so they move away from each other forming separate clusters. Once again, this metric shows in Figure 8 that the best tactic recovers faster than the Alignment strict tactic.
@@ 5 Related work
First notice that our work is not concerned with the modeling of actual pedestrian behaviours and trajectory planning [8, 9, 10]. Instead, our goal is to design simple tactics that lead to targeted features, like flocking.
@@
Likewise, our approach is different from usual protest models based on thresholds [11] or agent-based [12, 13, 14] approaches. Indeed, these works focus on how people decide to participate to a protest; they do not deal with protester mobility.
Models Network Lattice Flocking Fast
Random walk [15]
UTS/UXS [16, 17]
EAW [18]
SIP [19]
MIPS [20]
Boids [21]
Cucker-Smale [22]
Robot swarm [23, 24]
SVM [25, 26]
Our work
Table 1: Important models of collective walks found in literature that provide solutions to problems similar to ours. The specifies if they succeed at solving one of the issues we deal with.
5.1 Walks on networks
Instead, a number of works deal with flocking behaviors of walkers in a variety of situations. We summarize their relevant features in Table 1. It shows the limits of these works from our perspective. We discuss them in the following.
@@ 5.1 Walks on networks
In distributed systems, computer scientists proposed solutions to gathering problems where walkers follow a common distributed algorithm to meet on any connected graph [27, 28, 29]. In a sense, the main algorithms (UTS and UXS in Table 1[16, 17] from this community are similar tactics to our baseline, except they are deterministic.
It is important to note that labeling nodes, and letting walkers have the global knowledge to know all labels, are essential in their protocols. Moreover those studies only look for guaranties regarding gathering at long time. Unlike those, we had to ensure our walkers gather sufficiently fast and with no need for hard computations.
5.2 Flocking in free space
Flocking is a collective dynamic where groups of walkers move spontaneously in the same direction [22, 30]. Ranging from swarms to pedestrian crowds, numerous articles explored this behavior, in free space, since the Boids model [21]. The work of Vicsek [25] notably settled that alignment and noise were sufficient to produce flocking in the plan (SVM in Table 1). The combination of alignment and attraction is also an effective way to get gathering and flocking [31], as attraction avoids sprawling. Robot swarms also achieve flocking and similar objectives to ours [23]. However, that last example often uses the local communication model that exceeds our walkers capabilities.
In our case, trajectories of walkers are network constrained. As we are not in free space we could not, in a simple manner, define a notion of direction to describe walkers motion in a structure as loose as a graph (except for a lattice where you can define pairs of right/left, top/bottom direction). That is why we introduced instead the mobility score in Section 4.1.
5.3 Flocking and networks
Research on flocks and networks mostly do not study walks on graphs. They study the proximity graph of walkers: the graph of which walkers are in sight of another walker [32]. Researchers then study guaranties regarding the connectivity of this graph to evaluate if walkers achieve flocking.
However, a few articles explore flocking on simple networks. The motivation, in the context of statistical physics, is to use lattices as an approximation of free space to study active particles [20]. For example, in [26] the authors implement rules similar to ours to study the possible outcomes of their combinations on a line. They get results very similar to ours from Section 2.1. We also note that our baseline from Section 3.5 is very similar to the simple inclusion process (SIP in Table 1[19].
We can also cite algorithms that solve problems very close to ours on a network. There is [33] that studies the reciprocal problem of completely scattering walkers over a graph, while they are all gathered on a single node at the beginning. There is also the edge ant walk (EAW in Table 1[18], a bio-inspired model, where walkers communicate by leaving a trace of their passage, that can be used to patrol a graph.
Still, to our knowledge, we are the first to experiment rules for flocking that we can apply on any network. Such gap exists in the literature as such dynamic is not meaningful on many complex networks. Those often exhibit small-world and hub properties, making it very easy to gather fast when simply hopping from a high degree node to another at each step. Moreover, as phenomena in free spaces inspire flocking models, existing algorithms are then usually designed for application in free space and not on networks.
6 Conclusion
This paper proposes a framework for exploring what basic rules are crucial for walkers to gather and flock in a street network. Our findings indicate that the Alignment strict tactic is sufficient for flocking on a network. Alignment-based tactics are the most effective, and obtained groups are robust to break ups.
This framework is very general and may be extended in several ways. First, the city modelling may be improved with more urban data. For instance, one may use street width information to estimate the number of protesters needed to block each street. One may also use such data, as well as point-of-interest (POI) information, such as the presence of restaurants, stores, or theaters around protesters, to design more effective walking rules. Last but not least, this work opens the way to studying how obtained walks impact the underlying street network and traffic taking place on it.
Code and data.
We provide an implementation of our models in C, and a Python code result analysis: https://github.com/K-avi/protesting_on_graphs. We describe how to use it, including how to generate the obtained dataset, on the companion site: https://k-avi.github.io/protesting_on_graphs.
Acknowledgements
We warmly thank Ivan Mulot-Radojcic for his help with implementation and documentation. This work is funded in part by the ANR (French National Agency of Research) through the FiT-ANR-19-LCV1-0005 grant and by the CNRS through the MITI interdisciplinary programs.

### v2

# Fast Flocking of Protesters on Street Networks

@@ Abstract
We propose a simple model of protesters scattered throughout a city who want to gather into large and mobile groups. This model relies on random walkers on a street network that follow tactics built from a set of basic rules. Our goal is to identify the most important rules for fast flocking of walkers. We explore a wide set of tactics and show the central importance of a specific rule based on alignment. Other rules alone perform poorly, but our experiments show that combining alignment with them enhances flocking.
Keywords : street networks, gathering, flocking, tactic design, protests
@@ 1 Introduction
To do so, we model the city as a network of streets and intersections in Section 2. We then assume that protesters are biased random walkers on such a network. In Section 3, we present the set of basic rules that compose tactics walkers can use to move. We define, in Section 4, what flocking features groups must achieve and run extensive experiments to measure how effective each tactic is regarding those.
@@ 2.1 Street networks
The links of a street network generally represent street segments of very heterogeneous lengths [3]. Then, moves from a node to another one may have very different duration. In order to model this, we use a classical discretization procedure [4] that consists in splitting each link of the street network into pieces connected by evenly spaced nodes. In the obtained graph, each link represents a street slice of length close to a step $\delta$. Then, the walkers defined in previous section consistently make a move of length approximately $\delta$ at each time step.
@@ 2.2 Walkers
Given a network $G=(V,E)$, we consider a set $W$ of walkers numbered from $1$ to $|W|$. We denote the location of walker $i$ at time $t$ by $x_{i}(t)=v$, with $v\in V$. At each time step $t$, walker $i$ moves to node $x_{i}(t+1)\in N(v)$.
For any link $(u,v)$ in $E$, we also define the flux of walkers from $u$ to $v$ as $J_{u\rightarrow v}(t)=|\{i,x_{i}(t-1)=u,x_{i}(t)=v\}|$. We call group the set of walkers at a given node $v$ at a given time $t$: $g_{v}(t)=\{i,x_{i}(t)=v\}$. We denote by $n_{v}(t)=|g_{v}(t)|$ the number of walkers located at node $v$ at time $t$ and by $g(t)=|\{g_{v}(t),v\in V,g_{v}(t)\neq\emptyset\}|$ the number of non-empty groups at step $t$.
@@ 3.1 Available information
We assume that walkers have very limited capabilities: they have no access to the actual location of other walkers, even on neighbor nodes, and they have no long-term memory and no communication protocol. Instead, we assume they only have access to estimates of aggregated observables, such as the number of walkers on neighbor nodes and the flow intensity on surrounding links.
@@ 3.2 Criteria
  • Propulsion. The walker never goes back to its previous location $u$: the criterion has value 00 for $u$ and value $1$ for other neighbors of $v$.
@@ 3.4 Tactics
where $\alpha_{C}$ is the coefficient of criterion $C$, with $\sum_{C\in\mathcal{C}}\alpha_{C}=1$. Therefore, a tactic is defined by a set of criteria and their coefficients. We call strict tactic one that has all its coefficients, except one, set to zero and then always follows the same criterion.
@@ 3.5 Baseline
More formally, for each node $v$ at time $t$, a unique random neighbor $u$ of $v$ is chosen and all walkers $i$, such that $x_{i}(t)=v$, move to $u=x_{i}(t+1)$. This is equivalent to purely random walks of groups until they meet another group.
@@ 4.1 Flocking metrics
Definition 1 (cluster, gathering score $\rho(t)$).
@@
Definition 2 (mobility score $\mu(t)$).
The mobility $\mu_{i}(t)$ of walker $i$ is the number of distinct nodes a walker has already visited at time $t$: $\mu_{i}(t)=|\{v,\exists t^{\prime}\leq t,x_{i}(t^{\prime})=v\}|$. The mobility score $\mu(t)$ is the average walker mobility.
@@
Definition 3 (sprawling score $\sigma(t)$).
@@ 4.2 Extensive exploration of tactics
  • we consider $N$ walkers initially distributed uniformly at random on nodes,
@@
  • we set $\beta$ large enough to ensure each walking rule strictly follows its criterion,
@@
Figure 3: Plots showing the evolution of gathering, mobility and sprawling scores for the strict tactics and the best tactic. Gathering score is in log-log scale.
@@ 4.3 Best tactics
Finally, Figure 3 also displays the sprawling score for all considered tactics. We observe that the best tactic produces a greater sprawling score than Alignment. The sprawling, for those two tactics, is due to groups following each others when they detect another group on a neighbor node, without necessarily merging with it.
@@ 4.4 Interpretation
However, this rule alone is not perfectly effective at avoiding splitting. The walkers in the first group in a cluster will not always all choose the same node at an intersection, as illustrated in the two bottom pictures of Figure 4, as the different possible nodes all have a flux equal to 00. In this context, the Attraction rule allows walkers that split in the least chosen direction at an intersection to go backward and avoid loosing sight of the cluster.
@@ 5 Related work
Our approach is different from usual protest studies based on thresholds [7] or agent-based [8, 9, 10] model. Indeed, these works focus on how people decide to participate to a protest; they do not deal with protester mobility.
@@
In distributed systems, computer scientists proposed solutions to gathering problems where walkers follow a common distributed algorithm to meet on any connected graph [11, 12].
Flocking is a collective dynamic where groups of walkers move spontaneously in the same direction [13, 14, 15]. It has been largely studied in free space. Few articles explore flocking when trajectories of walkers are network constrained. For example, in [16] the authors implement rules similar to ours to study the possible outcomes of their combinations on a line. They get results very similar to ours in a single street.
Still, to our knowledge, we are the first to experiment rules for flocking that we can apply on any network.